I. A family buys a house worth $326,000. They pay $75,000 deposit and take a mortgage for the balance at J12=9% p.a. to be amortized over 30 years with monthly payments.
II. Fill out the loan amortization schedule provided in the solution template for the first 5 loan payments. What do you notice about the composition of the payment amount?
1
Expert's answer
2020-06-09T17:05:16-0400


I. 326 000-75 000=251 000

"2019.60\times360=727056"

interests=727 056-251 000=476056

"251 000=A\times\frac{1-(1.0075)^{-360}}{0.0075}"

A=2019.60

II. the loan amortization schedule:





As part of the amount of payments, the principal amount and interest for using the loan